Jill has $3.50 in nickels and dimes. If she has 50 coins, how many of each type of coin does she have?
Answer by fractalier(2101)

Let N be the number of nickels and D be dimes, so that
N + D = 50
Their values are 5N and 10D respectively, so that
5N + 10D = 350
Now let's multiply the first equation by 5 and subtract it from the second...we get
5N + 10D = 350
-(5N + 5D = 250)
so we have
5D = 100
D = 20
N = 30